OpenStudy (anonymous):

Simplify the expression 1 over 4 to the negative fourth power. A.256 B.16 C.negative 1 over 16 D.negative 1 over 256

OpenStudy (anonymous):

my answer is A.is it right

OpenStudy (zzr0ck3r):

\((\frac{1}{4})^{-4}=4^4\)

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

thanks

OpenStudy (zzr0ck3r):

\((\frac{a}{b})^{-c}=(\frac{b}{a})^c\)

OpenStudy (anonymous):

thanks for the help and the medal

OpenStudy (zzr0ck3r):

fo she z